an investor invested a total of $2,300 in two mutual funds. one fund earned 6% profit An investor invested a total of $2300 in two mutual funds. One fund earned a 6% profit, and the other a 2% profit. If the total profit was $74, how much was invested in each mutual fund?
Answer by stanbon(75887)

an investor invested a total of $2,300 in two mutual funds. one fund earned 6% profit An investor invested a total of $2300 in two mutual funds. One fund earned a 6% profit, and the other a 2% profit. If the total profit was $74, how much was invested in each mutual fund?
----------
Equation:
profit + profit = $74.00
--------------------
0.06x + 0.02(2300-x) = 74
-----
6x + 2*2300 - 2x = 7400
-----
4x = 7400-4600
4x = 2800
x = $700 (amt invested at 6%)
2300-x = $1600 (amt invested at 2%)
-----
Cheers,
Stan H.
--------